# Order-Optimal Global Convergence for Average Reward Reinforcement Learning via Actor-Critic Approach

### v2

@@ Abstract
This work analyzes average-reward reinforcement learning with general parametrization. Current state-of-the-art (SOTA) guarantees for this problem are either suboptimal or demand prior knowledge of the mixing time of the underlying Markov process, which is unavailable in most practical scenarios. We introduce a Multi-level Monte Carlo-based Natural Actor-Critic (MLMC-NAC) algorithm to address these issues. Our approach is the first to achieve a global convergence rate of $\tilde{\mathcal{O}}(1/\sqrt{T})$ without needing the knowledge of mixing time. It significantly surpasses the SOTA bound of $\tilde{\mathcal{O}}(T^{-1/4})$ where $T$ is the horizon length.
@@ 1 INTRODUCTION
Reinforcement Learning (RL) is a framework where an agent interacts with a Markovian environment and maximizes the reward it receives. The temporal dependence of the state transitions makes the problem of RL much more challenging than ordinary stochastic optimization, where data are selected in an independent and identically distributed (i.i.d.) manner. RL problems are typically analyzed via three setups: episodic, discounted reward with an infinite horizon, and average reward with an infinite horizon. The average reward framework is particularly significant for real-world applications, including robotics (Gonzalez et al., 2023), transportation (Al-Abbasi et al., 2019), communication networks (Agarwal et al., 2022), and healthcare (Tamboli et al., 2024).
One key approach to solving RL problems is model-based algorithms (Jaksch et al., 2010; Agrawal and Jia, 2017; Agarwal and Aggarwal, 2023) that directly learn the transition kernel from Markovian data. Unfortunately, such algorithms only work for small state spaces. Fueled by the practical need to learn policies for large state spaces, many model-free algorithms have emerged in the recent past where the policies are learned either from the $Q$-function or directly from the learned policy parameters. Some of these algorithms (Wei et al., 2020; Bai et al., 2024) use the observation that if the trajectory is divided into sub-trajectories that are mixing time apart, then these sub-trajectories become nearly independent. However, such approaches require the knowledge of mixing time which is not readily available in practice. Such an issue persists even if the MDP is assumed to be linear (Wei et al., 2020).
Recently, the authors of (Patel et al., 2024) showed that the algorithms that wait for orders of mixing time to have independent sub-trajectories require extremely large time horizons to even achieve sub-linear regret guarantees (even for mixing and hitting times of $10$, the time-horizon needed for sublinear guarantees is $6.6\times 10^{9}$). To alleviate these issues, (Patel et al., 2024) proposed an algorithm called Multi-level Actor Critic (MAC), where the estimates are obtained from Markovian trajectories utilizing the Multi-level Monte-Carlo (MLMC) gradient estimator (Giles, 2015) that does not require the knowledge of the mixing time. However, this approach is shown to achieve a global convergence rate of $\tilde{\operatorname*{\mathcal{O}}}(T^{-1/4})$ for general parameterized policies which is far away from the theoretical lower bound of $\Omega(T^{-1/2})$. This raises the following question.
@@
Algorithm Mixing Time Global Model-Free? Policy
Unknown? Convergence Parametrization
UCRL2 (Jaksch et al., 2010) - ${\tilde{\operatorname*{\mathcal{O}}}(1/\sqrt{T})}$ No Tabular
PSRL (Agrawal and Jia, 2017) - ${\tilde{\operatorname*{\mathcal{O}}}(1/\sqrt{T})}$ No Tabular
FOPO (Wei et al., 2021) No ${\tilde{\operatorname*{\mathcal{O}}}(1/\sqrt{T})}$ No Linear
OLSVI.FH (Wei et al., 2021) No $\tilde{\operatorname*{\mathcal{O}}}(1/T^{1/4})$ No Linear
MDP-EXP2 (Wei et al., 2021) No ${\tilde{\operatorname*{\mathcal{O}}}(1/\sqrt{T})}$ No Linear
MDP-OOMD (Wei et al., 2020) No ${\tilde{\operatorname*{\mathcal{O}}}(1/\sqrt{T})}$ Yes Tabular
PPGAE (Bai et al., 2024) No $\tilde{\operatorname*{\mathcal{O}}}(1/T^{1/4})$ Yes General
NAC-CFA (Wang et al., 2024) No $\tilde{\operatorname*{\mathcal{O}}}(1/T^{1/4})$ Yes General
MAC (Patel et al., 2024) Yes $\tilde{\operatorname*{\mathcal{O}}}(1/T^{1/4})$ Yes General
MLMC-NAC (Algorithm 1) Yes ${\tilde{\operatorname*{\mathcal{O}}}(1/\sqrt{T})}$ Yes General
Table 1: Summary of the key results on global convergence guarantees for average reward reinforcement learning.
@@
Is it possible to achieve a global convergence rate of $\tilde{\mathcal{O}}\left(T^{-1/2}\right)$ with general parameterized policies in the average-reward Markovian sampling setup without the algorithm requiring knowledge of the mixing time?
@@
Our Contribution: This article answers this question affirmatively. In particular, we introduce a Multi-level Monte Carlo-based Natural Actor-Critic (MLMC-NAC) algorithm that comprises two major components. The first component, referred to as the Natural Policy Gradient (NPG) subroutine, obtains an approximate NPG direction which is used to update the policy parameter. One sub-task of obtaining the NPG is estimating the advantage function. This is done via the second component, known as the Critic subroutine that achieves its desired target via Temporal Difference (TD) learning. Both NPG and critic subroutines apply MLMC-based gradient estimators that eliminate the use of the mixing time in the algorithm. We establish that MLMC-NAC achieves a global convergence rate of $\tilde{\operatorname*{\mathcal{O}}}(T^{-1/2})$ which is optimal in the order of the horizon length $T$ and far surpasses the state-of-the-art rate of $\tilde{\operatorname*{\mathcal{O}}}\left(T^{-1/4}\right)$.
The primary difficulty in handling Markovian trajectories, as opposed to independent trajectories (Liu et al., 2020; Mondal and Aggarwal, 2024), is that an accurate estimation of the policy gradient often requires the knowledge of mixing and hitting times (Bai et al., 2024). In most practical cases, these quantities are unknown. Moreover, the hitting time scales with the size of the state space, limiting the application of the underlying algorithm. As stated earlier, similar to (Patel et al., 2024), we employ an MLMC-based actor-critic structure that allows us to establish guarantees for large or infinite state spaces while eliminating the use of mixing time. However, the improvement in our convergence result yields from the following novel introductions.
  • We use NPG instead of the PG method used in (Patel et al., 2024). We prove that the global convergence error can be bounded by the sum of terms proportional to the bias and second-order error in NPG estimation (Lemma 1).
  • Since the critic updates underlie the NPG subroutine, the NPG estimation errors are inherently connected to critic estimation errors. In particular, we exhibit that NPG bias and second-order error are determined by their critic counterparts (Theorem 3).
  • Finally, we bound the critic errors in Theorem 4 that, combined with Theorem 3 and Lemma 1, yields our desired convergence rate.
  • We observe that both NPG and critic routines can be viewed as stochastic linear recursions where the stochasticity arises due to MLMC-based estimation. Theorem 2 analyses the convergence rate of such a general recursion which is utilized to prove Theorem 3 and 4. We would like to emphasize that it is due to the careful analysis of convergence bias in Theorem 2 that we get the optimal rate of $\tilde{\operatorname*{\mathcal{O}}}(T^{-1/2})$. Had we performed a naive convergence bias analysis of the general stochastic linear recursion in Theorem 2, the convergence rate of our algorithm would have been far away from being optimal.
@@
Related works: We will discuss the relevant works in three key areas as stated below. Our discussion primarily focuses on works that employ general parameterized policies. A summary of relevant works is available in Table 1.
  • Discounted Reward RL: We note that while optimal sample complexity has been studied for the discounted reward setups (Mondal and Aggarwal, 2024) for direct policy-gradient-based approaches, it assumes the availability of a simulator, enabling independent trajectory sampling. Markovian samplings are primarily studied via actor-critic approach (Xu et al., 2020; Gaur et al., 2024) which, unfortunately, have not achieved the optimal $\tilde{\operatorname*{\mathcal{O}}}(1/\epsilon^{2})$ sample complexity. Further, the analysis for discounted reward does not directly extend to the average reward setup since the sample complexity in the discounted case depends on $1/(1-\gamma)$, which blows up as $\gamma\to 1$.
  • Average Reward RL with direct Policy Gradient: The key work employing direct Policy Gradient for average reward setup with parametrized policy is (Bai et al., 2024). Unlike the discounted case, the policy gradient estimation is through the use of sub-trajectories that are large enough so that they are nearly independent. However, setting these trajectory lengths requires knowledge of the mixing time and the hitting time.
  • Average Reward RL with Actor-Critic approaches: To avoid creating independent subtrajectories, the policy gradient can be estimated by constructing value function and average reward estimates, updating based on observed samples. This is the basis of the actor-critic (AC) approach. Some studies on average reward reinforcement learning utilize an AC-based approach. For example, (Chen and Zhao, 2023) provides a local sample complexity of $\operatorname*{\mathcal{O}}(\epsilon^{-2})$ for vanilla AC, whereas (Panda and Bhatnagar, 2024) establishes a local sample complexity of $\mathcal{O}(\epsilon^{-2.08})$ for a Critic-Actor method. Another study, (Wang et al., 2024), introduces an AC method with a global sample complexity of $\mathcal{O}(\epsilon^{-4})$.
@@
    Actor-critic methods employing Multi-Level Monte Carlo (MLMC)-based estimation were first explored in (Suttle et al., 2023) and subsequently in (Patel et al., 2024). The former study demonstrates a local convergence, while the latter establishes a global convergence rate. Both studies investigate an MLMC-based AC algorithm combined with AdaGrad (Duchi et al., 2011), inspired by stochastic gradient descent (SGD) related work in (Dorfman and Levy, 2022). Unfortunately, none of these studies lead to an optimal global convergence rate which is the goal of our work.
@@ 2 SETUP
$d^{\pi_{\theta}}(s)=\lim_{T\rightarrow\infty}\operatorname{\mathbb{E}}_{\pi_{\theta}}\left[\frac{1}{T}\sum_{t=0}^{T}\mathbf{1}(s_{t}=s)\bigg{|}s_{0}\sim\rho\right]$ (1)
where $\operatorname{\mathbb{E}}_{\pi_{\theta}}$ denotes the expectation over the distribution of all $\pi_{\theta}$-induced trajectories and $\mathbf{1}(\cdot)$ is an indicator function. The above distribution also obeys $P^{\pi_{\theta}}d^{\pi_{\theta}}=d^{\pi_{\theta}}$. With this notation in place, we define the mixing time of an MDP.
@@
$t_{\mathrm{mix}}^{\theta}\coloneqq\min\left\{t\geq 1\bigg{|}\|(P^{\pi_{\theta}})^{t}(s,\cdot)-d^{\pi_{\theta}}\|\leq\dfrac{1}{4},\forall s\in\mathcal{S}\right\}$
We define $t_{\mathrm{mix}}\coloneqq\sup_{\theta\in\Theta}t^{\theta}_{\mathrm{mix}}$ as the the overall mixing time. In this paper, $t_{\mathrm{mix}}$ is finite due to ergodicity.
@@
$J^{\pi_{\theta}}:=\lim_{T\rightarrow\infty}\mathbb{E}_{\pi_{\theta}}\left[\frac{1}{T}\sum_{t=0}^{T}r(s_{t},a_{t})\bigg{|}s_{0}\sim\rho\right]$ (2)
@@
$Q^{\pi_{\theta}}(s,a)=$ $\mathbb{E}_{\pi_{\theta}}\Bigg{[}\sum_{t=0}^{\infty}\big{\{}r(s_{t},a_{t})-J(\theta)\big{\}}\bigg{|}s_{0}=s,a_{0}=a\Bigg{]}$
@@
$\nabla_{\theta}J(\theta)\!=\!\mathbb{E}_{(s,a)\sim\nu^{\pi_{\theta}}}\!\bigg{[}\!A^{\pi_{\theta}}(s,a)\nabla_{\theta}\log\pi_{\theta}(a|s)\!\bigg{]}$ (5)
@@
$F(\theta)=\operatorname{\mathbb{E}}_{(s,a)\sim\nu^{\pi_{\theta}}}\left[\nabla_{\theta}\log\pi_{\theta}(a|s)\otimes\nabla_{\theta}\log\pi_{\theta}(a|s)\right]$ (7)
@@
$L_{\nu^{\pi_{\theta}}}(\omega,\theta)\textstyle{=\dfrac{1}{2}\operatorname{\mathbb{E}}_{(s,a)\sim\nu^{\pi_{\theta}}}\left[\big{(}A^{\pi_{\theta}}(s,a)-\omega^{\top}\nabla_{\theta}\log\pi_{\theta}(a|s)\big{)}^{2}\right]}$ (8)
@@
$\nabla_{\omega}L_{\nu^{\pi_{\theta}}}(\omega,\theta)=F(\theta)\omega-\nabla_{\theta}J(\theta)$ (9)
@@
4:     $\triangleright$ Average reward and critic estimation
5:     for $h=0,1,\cdots,H-1$ do
6:         $s_{0}^{kh}\leftarrow s_{0}$, $Q_{kh}\sim\text{Geom}(1/2)$
7:         for $t=0,\dots,\max\{2^{Q_{kh}},T_{\max}\}-1$ do
@@
17:     $\triangleright$ Natural Policy Gradient (NPG) estimation
18:     for $h=H,H+1,\cdots,2H-1$ do
19:         $s_{0}^{kh}\leftarrow s_{0}$, $Q_{kh}\sim\text{Geom}(1/2)$
20:         for $t=0,\dots,\max\{2^{Q_{kh}},T_{\max}\}-1$ do
@@
29:     $\triangleright$ Policy update
30:     $\omega_{k}\leftarrow\omega^{k}_{2H}$, $\theta_{k+1}\leftarrow\theta_{k}+\alpha\omega_{k}$
31:end for
@@
$\nabla_{\eta}R(\theta,\eta)=\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\nu^{\pi_{\theta}}(s,a)\left\{\eta-r(s,a)\right\}$ (11)
@@
$\begin{split}&\nabla_{\zeta}E(\theta,\zeta)=\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\nu^{\pi_{\theta}}(s,a)\left\{\zeta^{\top}\phi(s)-Q^{\pi_{\theta}}(s,a)\right\}\phi(s)\end{split}$ (13)

### v3

# A Sharper Global Convergence Analysis for Average Reward Reinforcement Learning via an Actor-Critic Approach

@@ Abstract
This work examines average-reward reinforcement learning with general policy parametrization. Existing state-of-the-art (SOTA) guarantees for this problem are either suboptimal or hindered by several challenges, including poor scalability with respect to the size of the state-action space, high iteration complexity, and dependence on knowledge of mixing times and hitting times. To address these limitations, we propose a Multi-level Monte Carlo-based Natural Actor-Critic (MLMC-NAC) algorithm. Our work is the first to achieve a global convergence rate of $\tilde{\mathcal{O}}(1/\sqrt{T})$ for average-reward Markov Decision Processes (MDPs) (where $T$ is the horizon length), without requiring the knowledge of mixing and hitting times. Moreover, the convergence rate does not scale with the size of the state space, therefore even being applicable to infinite state spaces.
Machine Learning, ICML
@@ 1 Introduction
Reinforcement Learning (RL) is a framework where an agent interacts with a Markovian environment and maximizes the total reward it receives. The temporal dependence of the state transitions makes the problem of RL much more challenging than ordinary stochastic optimization, where data are selected in an independent and identically distributed (i.i.d.) manner. RL problems are typically analyzed via three setups: episodic, discounted reward with an infinite horizon, and average reward with an infinite horizon. The average reward framework is particularly significant for real-world applications, including robotics (Gonzalez et al., 2023), transportation (Al-Abbasi et al., 2019), communication networks (Agarwal et al., 2022), and healthcare (Tamboli et al., 2024). Model-based algorithms (Jaksch et al., 2010; Agrawal & Jia, 2017; Agarwal & Aggarwal, 2023) that learn the state transition kernel from Markovian trajectories, are well-known approaches for solving RL problems. However, these methods are typically limited to small state spaces. Policy Gradient (PG) methods, a cornerstone of RL, offer a model-free alternative that naturally supports function approximation (FA), making them well-suited for large state-action spaces. When the size of the state-action space, $SA$, is large or infinite, the framework of FA (also known as general parameterization) indexes the candidate policies by a $d$-dimensional parameter, $\theta$, where $d\ll SA$. Recently, some works have established global convergence guarantees for the average-reward setting with general policy parametrization, which we discuss below.
@@
Algorithm Infinite States Mixing Time Global Model-Free? Policy
and Actions? Unknown? Convergence Parametrization
UCRL2 (Jaksch et al., 2010) No -(1) ${\tilde{\operatorname*{\mathcal{O}}}(1/\sqrt{T})}$ No Tabular
PSRL (Agrawal & Jia, 2017) No -(1) ${\tilde{\operatorname*{\mathcal{O}}}(1/\sqrt{T})}$ No Tabular
MDP-OOMD (Wei et al., 2020)(2) No No ${\tilde{\operatorname*{\mathcal{O}}}(1/\sqrt{T})}$ Yes Tabular
PPGAE (Bai et al., 2024) No No $\tilde{\operatorname*{\mathcal{O}}}(1/T^{1/4})$ Yes General
PHPG (Ganesh et al., 2024) No No $\tilde{\operatorname*{\mathcal{O}}}(1/\sqrt{T})$ Yes General
NAC-CFA (Wang et al., 2024) Yes No $\tilde{\operatorname*{\mathcal{O}}}(1/T^{1/4})$ Yes General
MAC (Patel et al., 2024) Yes Yes $\tilde{\operatorname*{\mathcal{O}}}(1/T^{1/4})$ Yes General
MLMC-NAC (Algorithm 1) Yes Yes ${\tilde{\operatorname*{\mathcal{O}}}(1/\sqrt{T})}$ Yes General
Table 1: Summary of the key results on global convergence guarantees for average reward reinforcement learning. (1)Due to the model-based nature of the algorithms, these works use span of the optimal value function, $\mathrm{sp}(V^{*})$, (upper bounded by diameter) rather than mixing time. (2)This work also considers another algorithm that reduces mixing time to $\mathrm{sp}(V^{*})$ while achieving a higher rate of $\tilde{\mathcal{O}}(1/T^{1/3})$, though it still requires knowledge of $\mathrm{sp}(V^{*})$.
There are two main approaches in this context. The first is the direct PG method, where value functions are estimated directly from sampled trajectories (Bai et al., 2024; Ganesh et al., 2024). The second is the Temporal Difference (TD)-based policy evaluation approach, commonly known as the Actor-Critic (AC) method (Patel et al., 2024; Wang et al., 2024). Currently, the order-optimal $\tilde{\mathcal{O}}(T^{-1/2})$ convergence rate result, where $T$ is the horizon length, exists for direct PG method (Ganesh et al., 2024). However, these methods face several limitations, including poor state-action space scalability and reliance on knowledge of mixing and hitting times, which are unavailable in most practical scenarios. In contrast, existing AC algorithms circumvent these issues but are generally more challenging to analyze. More specifically, AC methods employ a TD-based critic to estimate the value function, which helps in reducing the variance. However, this reduction comes at the cost of introducing a bias, in addition to the inherent bias arising due to Markovian sampling. Direct PG methods using Markovian sampling leverage knowledge of mixing time to obtain nearly i.i.d. samples, whereas this would not suffice for AC methods due to the additional bias from the critic. As a result, the state-of-the-art bounds for AC is $\tilde{\operatorname*{\mathcal{O}}}(1/T^{1/4})$, which is suboptimal. This raises the following question:
@@
Is it possible to achieve a state-action space size independent global convergence rate of $\tilde{\mathcal{O}}\left(T^{-1/2}\right)$ for general parameterized policies in average reward infinite-horizon MDPs, using a practical algorithm that does not require the knowledge of mixing times?
@@
Our Contribution: This work answers this question affirmatively. In particular, we introduce a Multi-level Monte Carlo-based Natural Actor-Critic (MLMC-NAC) algorithm that comprises two major components. The first component, referred to as the Natural Policy Gradient (NPG) subroutine, obtains an approximate NPG direction which is used to update the policy parameter. One sub-task of obtaining the NPG is estimating the advantage function. This is done via the second component, known as the Critic subroutine that achieves its desired target via Temporal Difference (TD) learning. Both NPG and critic subroutines apply MLMC-based gradient estimators that eliminate the use of the mixing time in the algorithm. We establish that MLMC-NAC achieves a global convergence rate of $\tilde{\operatorname*{\mathcal{O}}}(T^{-1/2})$ which is optimal in the order of the horizon length $T$. The key contributions in this work are summarized as follows:
  • While existing AC analyses often use relatively loose bounds, we refine the analysis to achieve sharper results. Our first step towards this is to show that the global convergence error is bounded by terms proportional to the bias and second-order error in NPG estimation (Lemma 1). Since the critic updates underlie the NPG subroutine, the NPG estimation errors are inherently linked to critic estimation errors.
  • In prior AC works (Wang et al., 2024; Patel et al., 2024), the global convergence bound includes the critic error, $\operatorname{\mathbb{E}}\|\xi_{t}-\xi^{*}\|$, where $\xi_{t}$ is the critic estimate at time $t$ and $\xi^{*}$ is the true value. Instead, using Lemma 1 and Theorem 3, our analysis refines this term to $\|\operatorname{\mathbb{E}}[\xi_{t}]-\xi^{*}\|$, which can be significantly smaller than the previous estimate.
  • Bounding $\|\operatorname{\mathbb{E}}[\xi_{t}]-\xi^{*}\|$ still remains challenging due to Markovian noise. The critic update can be interpreted as a linear recursion with Markovian noise. Under i.i.d. noise, this term decays exponentially, but with Markovian noise, it can remain constant (Nagaraj et al., 2020). (Nagaraj et al., 2020) mitigates this by using one sample every $t_{\text{mix}}$ steps, requiring knowledge of the mixing time. Instead, we leverage MLMC to reduce the bias.
  • In Theorem 2, we establish a convergence rate for a generic stochastic linear recursion. Given that both the NPG and critic updates can be viewed as a stochastic linear update, this forms a basis for Theorems 3 and 4.
  • Theorem 1 proves the first $\tilde{\operatorname*{\mathcal{O}}}(T^{-1/2})$ global convergence result for AC methods.
@@
Related works: We will discuss the relevant works in two key areas as stated below. Our discussion primarily focuses on works that employ general parameterized policies. A summary of relevant works is available in Table 1.
Discussion on Practicality on direct PG methods: In direct PG methods, value function estimates are nearly unbiased but suffer from high variance, which scales with the size of the action space (Wei et al., 2020; Bai et al., 2024; Ganesh et al., 2024). Furthermore, the convergence results depend on the hitting time, which is at least the size of the state space, making the algorithm inapplicable to large or infinite state spaces. Finally, the implementation of these algorithms requires knowledge of mixing and hitting times, which is typically not available in practice and is difficult to estimate. In contrast, our algorithm does not require knowledge of mixing and hitting times.
Average Reward RL with Actor-Critic approaches: The authors of (Chen & Zhao, 2023; Panda & Bhatnagar, 2024; Suttle et al., 2023) provided local convergence guarantees for the AC based approaches. Recently, the global convergence for the AC methods in average reward setup have been studied in (Wang et al., 2024; Patel et al., 2024), where global sample complexity of $\tilde{\operatorname*{\mathcal{O}}}(T^{-1/4})$ is shown [^1] While (Wang et al., 2024) shows $\min_{1\leq t\leq T}(J^{*}-J(\theta_{t}))\leq\tilde{\operatorname*{\mathcal{O}}}(T^{-1/3})$, we note that the minimum error is strictly smaller than the average error notion that we consider in our work. Substituting the bound in Proposition 5 in their work in their average error decomposition in Step 1 of their Proof Outline yields an average error of $\tilde{\operatorname*{\mathcal{O}}}(T^{-1/4})$..
@@
We note that (Suttle et al., 2023; Patel et al., 2024) uses the Multi-Level Monte Carlo (MLMC)-based AC algorithm combined with AdaGrad (Duchi et al., 2011), inspired by stochastic gradient descent (SGD) related work in (Dorfman & Levy, 2022). Unfortunately, none of these studies lead to an optimal global convergence rate which is the goal of our work. We also note that the current state-of-the-art global convergence rate for AC methods in discounted MDPs is $\mathcal{O}(T^{-1/3})$ (Xu et al., 2020; Gaur et al., 2024), and the approaches proposed in this work have the potential to be applied in that setting.
@@ 2 Setup
where $\operatorname{\mathbb{E}}_{\pi_{\theta}}$ denotes the expectation over the distribution of all $\pi_{\theta}$-induced trajectories and $\mathbf{1}(\cdot)$ is an indicator function. The above distribution also obeys $(P^{\pi_{\theta}})^{\top}d^{\pi_{\theta}}=d^{\pi_{\theta}}$. With this notation in place, we define the mixing time of an MDP.
@@
where $\|\cdot\|_{\mathrm{TV}}$ denotes the total variation distance. We define $t_{\mathrm{mix}}\coloneqq\sup_{\theta\in\Theta}t^{\theta}_{\mathrm{mix}}$ as the the overall mixing time. This paper assumes $t_{\mathrm{mix}}$ to be finite.
@@
4:     for $h=0,1,\cdots,H-1$ do {Average reward and critic estimation}
5:        $s_{0}^{kh}\leftarrow s_{0}$, $Q_{kh}\sim\text{Geom}(1/2)$
6:        $\bar{Q}_{kh}\leftarrow 2^{Q_{kh}}$ if $2^{Q_{kh}}\leq T_{\max}$ else $\bar{Q}_{kh}\leftarrow 0$
7:        for $t=0,\dots,\bar{Q}_{kh}-1$ do
@@
17:     for $h=H,H+1,\cdots,2H-1$ do {Natural Policy Gradient (NPG) estimation}
18:        $s_{0}^{kh}\leftarrow s_{0}$, $Q_{kh}\sim\text{Geom}(1/2)$
19:        $\bar{Q}_{kh}\leftarrow 2^{Q_{kh}}$ if $2^{Q_{kh}}\leq T_{\max}$ else $\bar{Q}_{kh}\leftarrow 0$
20:        for $t=0,\dots,\bar{Q}_{kh}-1$ do
@@
29:     $\omega_{k}\leftarrow\omega^{k}_{2H}$, $\theta_{k+1}\leftarrow\theta_{k}+\alpha\omega_{k}$Policy update}
30:  end for
@@
$\begin{split}&\nabla_{\zeta}E(\theta,\zeta)\\ &=\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\nu^{\pi_{\theta}}(s,a)\left\{\zeta^{\top}\phi(s)-Q^{\pi_{\theta}}(s,a)\right\}\phi(s)\end{split}$ (14)